1. When 9.0 g of hydrogen gas, H2, reacts with oxygen gas, O2, 73.0 g of water is produced.    2H2 + O2 --> 2H2O a) What is the theoretical yield of water? b) What is the percentage yield?


2. Powdered zinc metal reacts with sulphur (S8) when heated to produce zinc sulphide.

a) What mass of product can be produced with 25 g of zinc?

b) What is the actual yield if the percentage yield was 80%?


3. The thermite reaction has been used to weld railroad rails, make bombs, and ignite solid rocket fuel. The equation of the reaction is:

a) What is the theoretical yield of iron if 100.0 g of Fe2O3 is used?

b) What is the percentage yield if 52.3 g of iron is actually produced?


1
Expert's answer
2022-01-18T09:27:01-0500

(1)

The equation for this reaction is

2H2 + O2 = 2H2O;

Having a mass of hydrogen we can find the quantity substance

n(H2) = m(H2)/M(H2);

M(H2) = 1 *2 = 2 g/mol

n(H2) = 9 g / 2 g/mol = 4.5 moles;

The amount of substance of hydrogen with the amount of substance of water as

a 2 to 2 or 1 to 1.

so n(H2O) = 4.5 moles

Theoretical mass of water will be

n(H2O) = m(H2O)/M(H2O);

M(H2O) = 1 *2 + 16 = 18 g/mol

m (H2O) 4.5 moles * 18 g/mol = 81 g.

so the theoretical yield of water is 81 g.

The percentage yield will be

m(H2O)practical / m(H2O)theoretical = 73.0 g / 81.0 g = 0.9012 or 90.12 %


(2)

The equation of the reaction is:

Zn+S =ZnS

Begin by determining the molar mass of each compound involved in the reaction. Using atomic masses from the periodic table, we will find the following:

M(ZnS) = 97.43 g/mol;

Ar(Zn) = 65.37 g/mol.

a) The mass of product determine from the equation of the reaction:

Zn + S= ZnS

m(ZnS)= "\frac{97.43 \times 25}{65.37}=37.26 g"

b) The actyal yield of ZnS is:

m(ZnS)=37.26×0.8=29.81 g. Answer:

  1. a) The theoretical mass of ZnS is 37.26 g;
  2. b) The actual yield of ZnS is: 29.81 g. 


(3)

The equation of the reaction is:

Fe2O3(s) + 2Al(s) → 2Fe(l) + Al2O3(s) ... balanced equation

Begin by determining the molar mass of each compound involved in the reaction.

Using atomic masses from the periodic table, we will find the following

"M(Fe_2O_3)=159.7g/mol;\\Ar(Fe)=55.85 g/mol."

a)

The theoretical yield of iron determine from the equation of the reaction:

"Fe_2O_3+2Al=2\space Fe +Al_2O_3\\m(Fe)=\frac{100 \times 2\times 55.85}{159.7}=69.9g"

b) The percentage yield of iron is:

n="\frac{52.3}{69.9}\times 100=74.8\%"

a) m(Fe)=69.9 g;

b) yield of iron is 74.8%